Grade 12Physical Chemistry

Calculate the energy associated with the first orbit of He+ . What is the radius of this orbit?

sudhanshu
12 Years agoGrade 12

1 Answer

Gitto Thomas Kuruthukulangarakoola
9 Years ago
Energy of orbit n, En = RH (Z2 / n2 ) and radius, rn = ao n2 / ZRH = 2.18 * 10-18 J; Z = 2; n = 1; ao = 52.9 pmE​n = ​2.18 * 10-18 J (4/1) = 8.72 * ​10-18 Jrn = 52.9 pm * 1 / 2 = 26.45 pm